Ticket Number: HEA-NOC/20090804-15 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20090804 16:36 UTC+1 Problem Start: 20090807 14:00 UTC+1
Ticket Update: 20090807 14:07 UTC+1 Problem End:
LYIT installing a new firewall
Possibility of brief outage to LYIT while routing is reconfigured.
20090804 16:27 OMG John O'Raw has requested a change of routing on Friday
afternoon to point at their new firewall. Opening ticket to track changes.
20090807 14:03 OMG Changed the static route.
This ticket can be monitored at http://www.hea.net/tickets/20090804-15
HEAnet Limited. Registered in Ireland, No. 275301.